# PoSh: Using Scene Graphs to Guide LLMs-as-a-Judge for Detailed Image Descriptions

### v1

@@ 5 Experiments
We extract sentence-level scene graphs using en_core_web_trf from Honnibal et al. (2020), a transformer trained to perform dependency parsing. To merge objects across these scene graphs while preserving attribute and relation attachments, we use maverick-mes-ontonotes from Martinelli et al. (2024) to perform co-reference resolution. Our QA scorer $\Psi$ is qwen-3-14b (Yang et al., 2025). We template evaluation questions for each scene graph component (as in Fig. 2), re-write candidate identifiers using $\Psi$ to improve fluency and then prompt $\Psi$ to answer each templated presence question by predicting a number between $1$ and $5$. We extract scores by taking the weighted average over the token logits for each number as in Liu et al. (2023). When determining object presence, we use a threshold of $2$, determined through tuning on a small hand-annotated validation set. We provide further implementation details and all prompts used in Section A.1.
@@
DOCENT is our new detailed description benchmark containing judgments from knowledgeable human annotators: granular mistake and omission spans for $300$ individual generations and coarse scaled rankings of mistakes, omissions and overall quality of $600$ paired generations. We evaluate granular metrics on this benchmark using macro F1 where we credit/penalize a model for predicting each annotated/unannotated token. Our coarse judgments are in the form $(text_{1},text_{2},score)$ where score indicates how much better or worse $text_{1}$ is than $text_{2}$. We evaluate each coarse metric $m$ by calculating its 1) pairwise accuracy (whether it picks the better text or a tie, using a tie threshold inferred from the gold tie rate) and 2) Spearman rank $\rho$ and Kendall’s $\tau$ correlations between $m(text_{1})-m(text_{2})$ and scorescore, a common practice in machine translation metric evaluation (Kocmi et al., 2021). More details can be found in Section A.3.3.
@@
CapArena (Cheng et al., 2025) contains $3,361$ images and $10,348$ detailed descriptions generated from $14$ current VLMs. $5,599$ pairs of these generations receive coarse judgments from human raters of the better generation (or “tie”). We include CapArena, which contains diverse images drawn from the web, to validate metric robustness. However, we note the dramatic simplicity of its images and references compared to those in DOCENT (see Table 1). $64\%$ of its images [^10] As measured by OneFormer (Jain et al., 2023) contain fewer than two objects and $95\%$ depict fewer than two people (compared to $27\%$ and $52\%$ in DOCENT). A metric is evaluated on CapArena at the caption-level (whether it picks the better text or a tie, using a tie threshold inferred from the gold tie rate) and at the model-level (Spearman’s rank and Kendall’s $\tau$ correlation between ELO rankings derived from metric predictions and gold judgments).
@@
Mistakes F1 $0.386$ $0.513$ $0.513$ $\mathbf{0.559}$
Omissions F1 $0.491$ $0.505$ $0.634$ $\mathbf{0.680}$
Granular Baselines
Our work is the first to introduce both a metric and a benchmark for granular evaluation of detailed descriptions. As such, this limits our baselines to those able to predict localized mistakes and omissions (i.e., the spans where errors occur). We consider two embedding-based approaches, using Qwen/Qwen3-Embedding-8B from Yang et al. (2025): 4GramEmbed, which embeds and compares $4$-grams from a generation and its reference, and SGEmbed, which embeds and compares components from the scene graphs of a generation and its reference. As these approaches (and PoSh) produce span scores, we report the maximum F1 scores for mistakes and omissions across all alerting thresholds. More details can be found in Section A.3.4.
Coarse Baselines
Though PoSh is a text-only reference-based metric, we select a representative set of reference-free (requiring only an image) and reference-based (requiring a gold standard) pointwise metrics (i.e. produce numerical scores) as our baselines. These include n-gram overlap metrics like BLEU (Papineni et al., 2002), ROUGE-L-Sum (See et al., 2017), METEOR (Banerjee & Lavie, 2005) and CIDER (Vedantam et al., 2015) and model-based metrics like SPICE (Anderson et al., 2016), CLIPScore (Hessel et al., 2021) and CAPTURE (Dong et al., 2024). Additionally, we consider several LLMs/VLMs-as-a-Judge [^11] CLAIR/Faithscore were not included due to complications with their codebases (Chan et al., 2023; Jing et al., 2024). Due to cost (estimated at \$$1,000$), we only evaluate DCScore (Ye et al., 2025) on DOCENT.: FLEUR (Lee et al., 2024), Prometheus (Kim et al., 2023), LLaVA-Critic (Xiong et al., 2025), DCScore Ye et al. (2025), Qwen-3 (Yang et al., 2025) and GPT4o/GPT5 in three settings (reference-free with image, reference-based without image and reference-based with image). More details can be found in Section A.3.5.
Reward Function
Finally, given the potential of a well-calibrated metric as a verifier in reinforcement learning (RL), we evaluate PoSh as a reward function. We train Qwen2.5-VL-7B on the $1,000$ images in DOCENT’s training set in two settings: 1) supervised fine-tuning (SFT), and 2) RL with DAPO (Yu et al., 2025) using PoSh. We collect coarse judgments (as in Section 4) for $40$ generation pairs from graduate students in NLP. More details can be found in Section A.3.6.
Table 3: Selected coarse metrics evaluated on DOCENT and CapArena, identified with $\Theta$ (parameter count, in billions), [Uncaptioned image] (requires a reference), [Uncaptioned image] (requires an image) and [Uncaptioned image] (replicable). “acc” indicates accuracy at predicting the better generation (or “tie”) in each judged pair. For DOCENT, $\rho$ / $\tau$ indicate the Spearman rank / Kendall’s $\tau$ correlations between differences in the metric and differences in the rank of the generations in each pair. For CapArena, $\rho$ / $\tau$ indicate the Spearman rank / Kendall’s $\tau$ correlations between model ELO rankings derived from metric scores and human judgments. Bold indicates the best replicable metric while underlining indicates the best metric overall. Gray cells indicate correlations that are not statistically significant at $\alpha=0.05$. PoSh beats nearly all baselines, including GPT4o, across both benchmarks in all settings (caption ranking of mistakes, omissions and overall quality & model ranking) while remaining replicable.
@@
LLaVa Critic 72 $\checkmark$ $\checkmark$ $\checkmark$ 62.8 0.412 0.351 57.0 0.509 0.430 66.8 0.546 0.461 64.0 0.987 0.934
@@
PoSh 14 $\checkmark$ $\checkmark$ 60.3 0.518 0.405 61.8 0.578 0.449 70.3 0.600 0.466 59.2 0.921 0.780
6 Results & Discussion
PoSh as a Granular Metric
Table 2 presents the performance of PoSh and our selected metrics on identifying the mistakes and omissions in DOCENT. Given the imbalanced nature of our data (where mistakes are infrequent and omissions are common), we report macro averages for each subtask, measuring how well each approach localizes errors within a generation and its reference respectively. First, we note that this task is difficult. The considerable room for improvement highlights the value of a benchmark like DOCENT that contains granular judgments of textual spans. Even so, PoSh achieves the highest F1 in mistake ($0.564$) and omission ($0.675$) localization. As its coarse scores are aggregated from these granular scores, this demonstrates its interpretability.
@@ 6 Results & Discussion
On DOCENT, across all three dimensions, PoSh outperforms every existing replicable metric (i.e., metrics not reliant on an API), yielding a $0.11$ increase in Spearman $\rho$ for mistakes ($26\%\uparrow$), a $0.07$ increase for omissions ($14\%\uparrow$) and a $0.05$ increase for overall quality ($10\%\uparrow$) over the next best. It even outperforms GPT4o (in all settings) and text-only GPT5 (on omissions and overall quality). Among all metrics, DCScore (Ye et al., 2025) proves best at predicting mistakes. However, its reliance on GPT4o to extract and verify factoids fails to achieve full coverage of reference detail, underperforming in predicting omissions and overall quality. Despite employing a smaller LLM, PoSh’s use of dependency parsing and coreference resolution to extract scene graphs avoids this.
On CapArena, PoSh achieves higher caption-level accuracies and model-ranking correlations than nearly every existing open-weight metric and GPT4o. The sole exception is LLaVa Critic, a much larger VLM-as-a-Judge (Xiong et al., 2025). This is driven in part by the simplicity of CapArena (see Table 1). On the subset of CapArena depicting three or more people ($167$ judgments), each of whom requires careful description, PoSh outperforms LLaVa Critic with model ranking correlations of $\mathbf{\rho=0.776,\tau=0.648}$ compared to $\mathbf{\rho=0.686,\tau=0.550}$. Thus, PoSh is robust to image type, excelling in visually complex cases that are of particular interest in detailed image description.
@@
In Table 7 of the Appendix, we report annotator agreement and aggregate preferences between SFT and DAPO with PoSh. In each dimension of interest, a PoSh generation earns a score between 2-2 and $2$ based on how much worse or better it is than its SFT counterpart. While PoSh-tuned generations had more mistakes (an average score of 0.243-0.243), these were incurred in service of much fewer missing details ($\mathbf{+0.432}$), resulting in higher overall quality ($\mathbf{+0.135}$). This speaks to the strength of PoSh when optimized directly.
DOCENT Leaderboard
Finally, in Fig. 9, we plot the PoSh scores of VLMs in describing the art in DOCENT. While closed models like Gemini 2.5 Pro lead, open models remain competitive. Improvements will require continued iteration, informed in part by PoSh’s granular scores.
@@ 8 Ethics Statement
The judgments in DOCENT were collected under an approved IRB protocol with all annotator data anonymized and participants receiving fair compensation (at $\mathdollar 22$/hour) for their time and expertise.
@@ 9 Reproducibility Statement
    publishing the code for both our metric and our metric evaluations at https://github.com/amith-ananthram/posh
@@ A.1 PoSh
A.1.1 Scene Graph Extraction
@@ A.2.1 Generations
    llava-v1.6-mistral-7b-hf on HuggingFace (Liu et al., 2024)
@@
    Molmo-7B-D-0924 on HuggingFace (Deitke et al., 2024)
@@
We use the same prompt (included below). For LLaVA-1.5-7B and Molmo-D-7B, we use nucleus sampling Holtzman et al. (2019) with $p=0.9$ and a temperature of $0.7$.
@@
Figure 4: The beginning of our granular annotation instructions. The full instructions can be found on our GitHub: https://anonymous.4open.science/r/posh/docent/annotation_instructions/granular.pdf.
Figure 5: The beginning of our coarse annotation instructions. The full instructions can be found on our GitHub: https://anonymous.4open.science/r/posh/docent/annotation_instructions/coarse.pdf.
@@
A.2.2 Avoiding Double Penalties
In Kasai et al. (2022), after identifying an error in precision, the authors correct the error before annotating recall. This avoids doubly penalizing a description for errors in specificity which would unfairly favor more generic descriptions (which are only penalized once, for recall). We instruct our annotators to do the same. Below, we have manually identified a few cases of its application from our granular annotations though we note that annotators found our models to be reasonably precise (Table 5) so this rule was infrequently required.
Due to the length of the generations and descriptions in DOCENT, please consult our codebase for example judgments: https://anonymous.4open.science/r/posh/docent

### v2

@@ 5 Experiments
We extract sentence-level scene graphs using en_core_web_trf from spacy, a transformer trained to perform dependency parsing. To merge objects across these scene graphs while preserving attribute and relation attachments, we use maverick-mes-ontonotes from martinelli-etal-2024-maverick to perform co-reference resolution. Our QA scorer $\Psi$ is qwen-3-14b (yang2025qwen3). We template evaluation questions for each scene graph component (as in Fig. 2), re-write candidate identifiers using $\Psi$ to improve fluency and then prompt $\Psi$ to answer each templated presence question by predicting a number between $1$ and $5$. We extract scores by taking the weighted average over the token logits for each number as in liu2023g. When determining object presence, we use a threshold of $2$, determined through tuning on a small hand-annotated validation set. We provide further implementation details and all prompts used in Section A.1.
@@
DOCENT is our new detailed description benchmark containing judgments from knowledgeable human annotators: granular mistake and omission spans for $300$ individual generations and coarse scaled rankings of mistakes, omissions and overall quality of $600$ paired generations. We evaluate granular metrics on this benchmark using macro F1 where we credit/penalize a model for predicting each annotated/unannotated token. Our coarse judgments are in the form $(text_{1},text_{2},score)$ where score indicates how much better or worse $text_{1}$ is than $text_{2}$. We evaluate each coarse metric $m$ by calculating its 1) pairwise accuracy (whether it picks the better text or a tie, using a tie threshold inferred from the gold tie rate) and 2) Spearman rank $\rho$ and Kendall’s $\tau$ correlations between $m(text_{1})-m(text_{2})$ and scorescore, a common practice in machine translation metric evaluation (kocmi2021ship). More details can be found in Section A.3.1.
@@
CapArena (cheng2025caparena) contains $3,361$ images and $10,348$ detailed descriptions generated from $14$ current VLMs. $5,599$ pairs of these generations receive coarse judgments from human raters of the better generation (or “tie”). We include CapArena, which contains diverse images drawn from the web, to validate metric robustness. However, we note the dramatic simplicity of its images and references compared to those in DOCENT (see Table 1). $64\%$ of its images [^10] As measured by OneFormer (jain2023oneformer) contain fewer than two objects and $95\%$ depict fewer than two people (compared to $27\%$ and $52\%$ in DOCENT). A metric is evaluated on CapArena at the caption-level (whether it picks the better text or a tie, using a tie threshold inferred from the gold tie rate) and at the model-level (Spearman’s rank and Kendall’s $\tau$ correlation between ELO rankings derived from metric predictions and gold judgments).
Granular Baselines
Our work is the first to introduce both a metric and a benchmark for granular evaluation of detailed descriptions. As such, this limits our baselines to those able to predict localized mistakes and omissions (i.e., the spans where errors occur). We consider two embedding-based approaches, using Qwen/Qwen3-Embedding-8B from yang2025qwen3: 4GramEmbed, which embeds and compares $4$-grams from a generation and its reference, and SGEmbed, which embeds and compares components from the scene graphs of a generation and its reference. As these approaches (and PoSh) produce span scores, we report the maximum F1 scores for mistakes and omissions across all alerting thresholds. More details can be found in Section A.3.3.
Coarse Baselines
Though PoSh is a text-only reference-based metric, we select a representative set of reference-free (requiring only an image) and reference-based (requiring a gold standard) pointwise metrics (i.e. produce numerical scores) as our baselines. These include n-gram overlap metrics like BLEU (papineni2002bleu), ROUGE-L-Sum (see2017get), METEOR (banerjee2005meteor) and CIDER (vedantam2015cider) and model-based metrics like SPICE (anderson2016spice), CLIPScore (hessel2021clipscore) and CAPTURE (dong2024benchmarking). Additionally, we consider several LLMs/VLMs-as-a-Judge [^11] CLAIR/Faithscore were not included due to complications with their codebases (chan2023clair; jing2024faithscore). Due to cost (estimated at \$$1,000$), we only evaluate DCScore (ye2025painting) on DOCENT.: FLEUR (lee2024fleur), Prometheus (kim2023prometheus), LLaVA-Critic (xiong2024llavacritic), DCScore ye2025painting, Qwen-3 (yang2025qwen3) and GPT4o/GPT5 in three settings (reference-free with image, reference-based without image and reference-based with image). More details can be found in Section A.3.4.
Reward Function
Finally, given the potential of a well-calibrated metric as a verifier in reinforcement learning (RL), we evaluate PoSh as a reward function. We train Qwen2.5-VL-7B on the $1,000$ images in DOCENT’s training set in two settings: 1) supervised fine-tuning (SFT), and 2) RL with DAPO (yu2025dapo) using PoSh. We collect coarse judgments (as in Section 4) for $40$ generation pairs from graduate students in NLP. More details can be found in Section A.3.5.
6 Results & Discussion
@@
Mistakes F1 $0.503$ $0.483$ $0.514$ $\mathbf{0.580}$
Omissions F1 $0.499$ $0.641$ $0.658$ $\mathbf{0.680}$
PoSh as a Granular Metric
Table 2 presents the performance of PoSh and our selected metrics on identifying the mistakes and omissions in DOCENT. Given the imbalanced nature of our data (where mistakes are infrequent and omissions are common), we report macro averages for each subtask, measuring how well each approach localizes errors within a generation and its reference respectively. First, we note that this task is difficult. The considerable room for improvement highlights the value of a benchmark like DOCENT that contains granular judgments of textual spans. Even so, PoSh achieves the highest F1 in mistake ($0.580$) and omission ($0.680$) localization. As its coarse scores are aggregated from these granular scores, this demonstrates its interpretability.
Table 3: Selected coarse metrics evaluated on DOCENT and CapArena, identified with $\Theta$ (parameter count, in billions), [Uncaptioned image] (requires a reference), [Uncaptioned image] (requires an image) and [Uncaptioned image] (replicable). “acc” indicates accuracy at predicting the better generation (or “tie”) in each judged pair. For DOCENT, $\rho$ / $\tau$ indicate the Spearman rank / Kendall’s $\tau$ correlations between differences in the metric and differences in the rank of the generations in each pair. For CapArena, $\rho$ / $\tau$ indicate the Spearman rank / Kendall’s $\tau$ correlations between model ELO rankings derived from metric scores and human judgments. Bold indicates the best replicable metric while underlining indicates the best metric overall. Gray cells indicate correlations that are not statistically significant at $\alpha=0.05$. PoSh beats nearly all baselines, including GPT4o, across both benchmarks in all settings (caption ranking of mistakes, omissions and overall quality & model ranking) while remaining perfectly replicable.
@@
LLaVa Critic $72$ $\checkmark$ $\checkmark$ $\checkmark$ $\mathbf{\underline{62.8}}$ $0.412$ $0.351$ $57.0$ $0.509$ $0.430$ $66.8$ $0.546$ $0.461$ $\mathbf{\underline{64.0}}$ $\mathbf{\underline{0.987}}$ $\mathbf{\underline{0.934}}$
DCScore $\checkmark$ $\checkmark$ $\underline{62.8}$ $\underline{0.541}$ $0.422$ $54.0$ $0.395$ $0.298$ $62.8$ $0.471$ $0.362$ - - -
@@
GPT5 $\checkmark$ $62.5$ $0.511$ $\underline{0.423}$ $53.2$ $0.421$ $0.332$ $68.0$ $0.540$ $0.440$ $59.1$ $0.956$ $0.846$
PoSh $14$ $\checkmark$ $\checkmark$ $60.7$ $\mathbf{0.519}$ $\mathbf{0.405}$ $\mathbf{\underline{62.7}}$ $\mathbf{\underline{0.581}}$ $\mathbf{\underline{0.451}}$ $\mathbf{\underline{70.7}}$ $\mathbf{\underline{0.599}}$ $\mathbf{\underline{0.466}}$ $59.2$ $0.931$ $0.796$
@@
On DOCENT, across all three dimensions, PoSh outperforms every existing replicable metric (i.e., metrics not reliant on an API), yielding a $0.11$ increase in Spearman $\rho$ for mistakes ($25\%\uparrow$), a $0.07$ increase for omissions ($14\%\uparrow$) and a $0.05$ increase for overall quality ($9\%\uparrow$) over the next best. It even outperforms GPT4o (in all settings) and text-only GPT5 (on omissions and overall quality). Among all metrics, DCScore (ye2025painting) proves best at predicting mistakes. However, its reliance on GPT4o to extract and verify factoids fails to achieve full coverage of reference detail, underperforming in predicting omissions and overall quality. Despite employing a smaller LLM, PoSh’s use of dependency parsing and coreference resolution to extract scene graphs avoids this.
On CapArena, PoSh achieves higher caption-level accuracies and model-ranking correlations than nearly every existing open-weight metric and GPT4o. The sole exception is LLaVa Critic, a much larger VLM-as-a-Judge (xiong2024llavacritic). This is driven in part by the simplicity of CapArena (see Table 1). On the subset of CapArena depicting three or more people (167 judgments), each of whom requires careful description, PoSh outperforms LLaVa Critic with model ranking correlations of $\mathbf{\rho=0.727,\tau=0.581}$ compared to $\mathbf{\rho=0.686,\tau=0.550}$. Thus, PoSh is robust to image type, excelling in visually complex cases that are of particular interest in detailed image description.
@@
In Table 7 of the Appendix, we report annotator agreement and aggregate preferences between SFT and DAPO with PoSh. In each dimension of interest, a PoSh-tuned generation earns a score between 2-2 and $2$ based on how much worse or better it is than its SFT counterpart. While PoSh-tuned generations had more mistakes (an average score of 0.243-0.243), these were incurred in service of much fewer missing details ($\mathbf{+0.432}$), resulting in higher overall quality ($\mathbf{+0.135}$). This speaks to the strength of PoSh when optimized directly. Given recent progress in generating synthetic detailed descriptions (li2024densefusion), PoSh-tuning could be freely scaled in post-training. Moreover, as PoSh produces localized granular scores, it supports token-level guidance (yang2023preference), an exciting direction to explore in future work.
PoSh Subcomponent Evaluation
PoSh relies on two subcomponents, scene graph extraction and scene graph element verification through question answering. As downstream errors may propagate from these components, we validate each through comparison against hand annotated examples from DOCENT. We find that PoSh’s scene graphs are high quality, with average element F1 of $0.892$. Similarly, in verifying these scene graph elements, PoSh exhibits strong alignment with human raters, with an F1 of $0.852$. We provide additional details and analysis in Section A.4.2.
PoSh Runtime
A core enabler of PoSh’s performance and interpretability is its thorough granular evaluation. It achieves this efficiently through inference optimizations like continuous batching and prefix caching (kwon2023efficient). PoSh scores the $400$ examples in DOCENT in $15$ minutes, or one every $2$ seconds, on a single H100 GPU. In contrast, DCScore (ye2025painting) takes upwards of $2$ hours due to its heavy use of GPT4 (one description every $25$ seconds). As manual evaluation takes $18$ minutes per description (see Table 1), PoSh effectively balances quality and cost.
7 DOCENT Leaderboard
Finally, in Fig. 9, we plot the PoSh scores of VLMs in describing the art in DOCENT. While closed models like Gemini 2.5 Pro lead, open models remain competitive. Improvements will require continued iteration, informed in part by insights gained from analyzing PoSh’s granular scores.
@@ 8 Conclusion
9 Limitations
Recent efforts have explored using structural priors to guide generation (e.g. wang2025llava). As these methods extract and describe structure from images, and as PoSh extracts and validates structure from text, we do not expect PoSh to be biased towards them in an evaluation setting. Nevertheless, as these models become publicly available, this requires experimental validation.
@@ 10 Ethics Statement
The judgments in DOCENT were collected under IRB protocol AAAV6216 with all annotator data anonymized and participants receiving fair compensation (at $\mathdollar 22$/hour) for their time and expertise.
@@ 11 Reproducibility Statement
    publishing the code for both our metric and our metric evaluations at https://github.com/amith-ananthram/posh; this implementation supports batch invariance, ensuring perfect reproducibility of our results on an H100 GPU with CUDA 12.8
@@ A.1 PoSh
A.1.1 Comparison to Davidsonian Scene Graph
There are several differences between the Davidsonian Scene Graph (DSG) metric (chodavidsonian) and PoSh. While the emphasis of DSG is on evaluating text-to-image models, PoSh was designed specifically for detailed image descriptions and is tailored to the unique challenges they pose.
First, the emphasis of DSG is on evaluating text-to-image models: it compares an image to scene graph elements extracted from a prompt text through visual question answering. As such, it cannot serve as a reference-based metric for evaluating detailed image descriptions. Allowing the use of references is important. Downstream tasks have context-specific requirements that can only be specified with references. This is especially true in accessibility (deutsch2022limitations).
Additionally, DSG was designed for image generation prompts at most three sentences long (in contrast, PoSh is able to compare generations and references that are $10$ - $20$ sentences long); DSG prompts GPT-3.5 for atomic propositions that are not localized to text spans (in contrast, PoSh grounds its coarse scores in localized granular scores with open models, allowing error visualization, better interpretability and replicability), DSG’s atomic propositions do not have special handling for entity collisions (in contrast, PoSh tests discriminating identifiers for colliding entities to allow unique validation of their presence), and finally, DSG applied to detailed image descriptions measures only precision, validating the presence of generation elements in its source image (in contrast, PoSh measures both precision and recall, penalizing generations for omitting important details).
These differences in design and purpose become clear when evaluating DSG against the judgments in PoSh, where its accuracies are near chance and its correlations are near zero (see Table 6).
A.1.2 Scene Graph Extraction
@@ A.2.1 Generations
    llava-v1.6-mistral-7b-hf on HuggingFace (liu2024improved)
@@
    Molmo-7B-D-0924 on HuggingFace (deitke2024molmo)
@@
We use the same prompt (included below). For LLaVA-1.5-7B and Molmo-D-7B, we use nucleus sampling holtzmancurious with $p=0.9$ and a temperature of $0.7$.
@@
A.2.2 Avoiding Double Penalties
In kasai-etal-2022-transparent, after identifying an error in precision, the authors correct the error before annotating recall. This avoids doubly penalizing a description for errors in specificity which would unfairly favor more generic descriptions (which are only penalized once, for recall). We instruct our annotators to do the same.
Due to the length of the generations and descriptions in DOCENT, please consult our codebase for example judgments: https://github.com/amith-ananthram/posh/tree/main/docent/examples/granular
Figure 4: The beginning of our granular annotation instructions, also hosted on our GitHub.
Figure 5: The beginning of our coarse annotation instructions, also hosted on our GitHub.